Which of the following 0.1M aqueous solutions has the lowest pH at 25 C?

A. K2HPO3

B. K2PO3

C. NaH2PO3

D. Li3PO3

E. NaOH

(the answer is C, but can someone explain why)
Answer #1

lowest pH --> will be the most acidic

then

compare

A. K2HPO3 = HPO3-2 forms, only 1 proton to donate

B. K2PO3 = PO4-3 will have no protons

C. NaH2PO3 = H2PO3- has 2 protons to donate,

D. Li3PO3 = PO4-3 has no protons to donate

E. NaOH = this is a strong base

then, we need the most acidic

acid = proton donor

the species with most protons

NaH2PO4